# Efficient and Workload-Aware LLM Serving via Runtime Layer Swapping and KV Cache Resizing

### v1

@@ 3.4 KVResizer: Elastic KVC Resizing
To support bursty workloads and fluctuating memory demands, MorphServe integrates dynamic layer swapping with KVResizer, a mechanism for elastic resizing of key-value cache (KVC) blocks. This section addresses two key questions: (1) how KVResizer dynamically allocates and releases KVC blocks in response to runtime memory pressure, and (2) how it collaborates with layer swapping to maintain serving efficiency under peak load.
@@
KVResizer extends PagedAttention [33] with kernel-level support for on-demand KVC block allocation/deallocation, implemented through memory mapping without requiring kernel recompilation. All resizing operations are executed asynchronously using separate CUDA streams to avoid interference with ongoing decoding.
Unlike static preallocation strategies (e.g., in vLLM [33]), KVResizer adjusts KVC capacity dynamically based on real-time memory availability. Once the pressure subsides, both temporary KVC blocks and quantized layers are released and restored to their full-precision state, ensuring memory reuse and accuracy recovery.
@@ 4 Experiment
Figure 4: MorphServe provides the best latency–accuracy tradeoff across four models and two traces, with four datasets. MorphServe in accuracy mode (dark green) reduces P95 TTFT by $\times2$3.9×3.9$\times$3.9 × compared to full-precision serving while maintaining comparable generation quality. In performance mode (light green), MorphServe consistently outperforms INT4 quantized models in output quality with no additional latency overhead.
Evaluation Setup. We evaluate MorphServe across a diverse range of LLM architectures, workload traces, and tasks. We consider four representative models: Vicuna 7B v1.5 [42], Llama 2 7B [64], Llama 3 8B [22], and CodeLlama 34B [56], spanning multiple scales and attention types—including Multi-Head Attention (MHA) [65] and Grouped-Query Attention (GQA) [3]. We test two real-world LLM inference workload traces: the BurstGPT trace [67] and the Azure LLM Inference trace [61, 4]. We report results from a representative 72-second trace snippet (Figure 1) for both workloads, though MorphServe is effective across the full traces. The request arrival rates of each trace are downscaled by 1.75$\times$ and 4.75$\times$ to fit our hardware environment. To evaluate generation quality, we use four public datasets: GovReport [27] and Multi-News [15] (long-form summarization), QMSum [77] (query-based summarization), and DuReader [23] (reading comprehension). For each test, we align workload timestamps with context passages from the datasets. Prompt and response lengths are set to 512 and 256 tokens for Vicuna 7B v1.5 and Llama 2 7B, and to 1024 and 512 tokens for Llama 3 8B and CodeLlama 34B. We report F1 and Rouge-L scores to assess generation quality. End-to-end experiments for Vicuna 7B v1.5, Llama 2 7B, and Llama 3 8B are conducted on an NVIDIA L4 GPU with 24 GB HBM and 256 GB of CPU DRAM, while CodeLLaMA 34B is evaluated on an A100 server with 80 GB HBM and 2 TB of CPU DRAM.
@@
Implementation. MorphServe is implemented on top of SwiftLLM [31, 58], a lightweight and modular LLM inference framework that reproduces vLLM [33] performance with simplified components. We added approximately 2,200 lines of Python and 500 lines of C++/CUDA to support MorphServe’s optimized KVC management and attention kernel extensions, which enable efficient layer swapping and KVC resizing at runtime.
Baselines. We include the full-precision (FP16) model as an upper-bound reference and an INT4 quantized model as a static compression baseline. For quantization, we adopt AWQ [38] due to its efficient inference kernel support; however, the setup is compatible with any post-training quantization method and can be replaced accordingly. To evaluate the flexibility of MorphServe, we configure it in two runtime modes: In accuracy mode, MorphServe prioritizes output quality by raising the threshold for triggering layer swapping and limiting the number of quantized layer replacements, thereby minimizing accuracy degradation. In performance mode, MorphServe enables more aggressive layer swapping to improve throughput and reduce latency under memory pressure. All baselines and MorphServe configurations are evaluated on the same serving engine to ensure fair comparison.
@@ 4.1 Main Results
TTFT and Accuracy. As shown in Figure 4, MorphServe significantly reduces P95 TTFT latency while preserving output quality in all model-trace-dataset configurations. Compared to full-precision baselines, MorphServe reduces the P95 TTFT by 2.9×2.9$\times$2.9 ×15.7×15.7$\times$15.7 × ($\times2$3.9×3.9$\times$3.9 × in accuracy mode and 3.4×3.4$\times$3.4 ×19.5×19.5$\times$19.5 × in performance mode) while maintaining quality within $0.51\%$$3.82\%$ degradation on F1 or Rouge-L scores, as low as $0.11\%$$2.18\%$ in accuracy mode. In contrast, static quantization exhibits $2.34\%$-$9.47\%$ degradation compared to full-precision inference, due to suffering from persistent quality loss across the entire serving lifetime. In particular, MorphServe excels in long-context datasets such as GovReport, leveraging LayerSwapper and KVResizer to optimize memory and computing efficiency. MorphServe with different configurations (green stars) visualizes the ability to navigate the latency-accuracy Pareto frontier, offering the best balance of performance and quality based on real-time workload shifts.
@@
Workload Adaptation and Saturation Resilience. As shown in Figure 5, MorphServe adaptively manages KVC block capacity in response to fluctuating load. In the full-precision baseline, KVC usage saturates the static capacity limit during peak periods, resulting in elevated queueing delays, request preemption, and frequent KVC swapping, which can lead to SLO violations. Static quantization, while reducing the memory footprint, degrades model accuracy and underutilizes GPU memory, even during low-load periods. MorphServe attaches new blocks during bursty traffic and releases them as load subsides, enabled by the synergistic LayerSwapper and KVResizer mechanism. MorphServe improves overall KVC memory utilization and output accuracy by $29.29\%$ and $3.58\%$, respectively, compared to static quantization. The adaptation allows MorphServe to expand KVC usage by up to $32.97\%$ beyond the full-precision limit when needed, and reduce the queueing delay by up to 3.8×3.8$\times$3.8 ×. MorphServe also mitigates request preemption and KVC swapping under saturation conditions. This enhances system responsiveness and improves token-level efficiency, contributing to reduced TPOT and end-to-end request latency.
Figure 6: MorphServe delays saturation and achieves up to 1.83×1.83$\times$1.83 × throughput than full-precision serving under increasing request rates.
@@
Throughput. In Figure 6, we compare MorphServe with baselines on DuReader under varying request rates. All configurations maintain low TTFT at low RPS, but as load increases, full-precision inference encounters the threshold, where TTFT spikes abruptly due to memory exhaustion and queueing delays. In contrast, MorphServe consistently pushes back this saturation point, achieving 1.6×1.6$\times$1.6 ×1.83×1.83$\times$1.83 × higher throughput than full-precision serving across all evaluated models.
Figure 7: MorphServe incurs negligible runtime overhead while improving tail TPOT latency. MorphServe (green) achieves comparable average TPOT latency to the full-precision baseline (blue), while reducing P99 latency by up to 1.23×1.23$\times$1.23 ×. Performance mode (light green) improves the average TPOT by up to 1.17×1.17$\times$1.17 × through aggressive layer morphing.
CPU Memory Overhead. Compared to full-precision, MorphServe introduces modest additional host memory usage by maintaining a mixed set of full-precision and quantized variants of transformer layers. Fortunately, the overhead is bounded: quantized weights (e.g., W8, W4, W3) are significantly smaller than their full-precision counterparts, and the combined memory footprint typically does not exceed 2$\times$ the original model size. Moreover, multi-GPU deployment of large models or model replicas hosted on multiple GPUs within the same node can share a single copy of quantized, CPU-memory-residential model weights across GPUs, eliminating redundant CPU memory consumption. In our experiments on an NVIDIA A100 8$\times$ 80GB server with 2 TB of host memory, the total memory footprint, including both swapped-out full-precision layers and INT4 quantized variants of CodeLlama 34B, accounted for only $4.42\%$ of available host memory, introducing negligible memory bandwidth and capacity overhead. These results confirm that MorphServe’s host memory footprint is practical and sustainable for both cloud-scale and high-end edge deployments.
@@
Runtime Performance Overhead. Figure 7 presents the cumulative distribution (CDF) of time-per-output-token (TPOT) across two datasets and four models under the Azure LLM trace. MorphServe delivers average TPOT comparable to full-precision serving while improving P95 and P99 TPOT tail latency by up to 1.06×1.06$\times$1.06 × and 1.23×1.23$\times$1.23 ×, respectively. These gains are achieved by eliminating request preemption stalls and avoiding KVC swapping or recomputation—two primary sources of long-tail delays. The performance mode of MorphServe reduces average TPOT by $1\times1$-1.17×1.17$\times$1.17 ×, while the accuracy mode introduces overhead of up to 1.06×1.06$\times$1.06 × as it preserves more full-precision layers and applies stricter thresholds for layer swapping. In accuracy mode, this conservative strategy increases memory usage and may lead to occasional queueing delays under load. The TPOT gain from MorphServe is due to faster inference on quantized layers, and the highly efficient kernels on layer swapping (e.g., $\sim$6 ms for a Llama 2 7B INT4 attention layer). These results confirm that MorphServe introduces negligible runtime overhead while effectively reducing tail latency.
MorphServe supports an optional offline calibration step to pre-compute layer sensitivity scores for more accurate morphing decisions. While this process improves accuracy-latency tradeoffs, it is not required for MorphServe to function. For a model, the calibration is a once-for-all process. Given its offline nature and minimal duration, the overhead is negligible and acceptable in practice. Details of the calibration procedure and associated cost are provided in the Appendix.
@@ 5 Conclusion
This paper presents MorphServe, a novel workload-aware LLM serving framework based on morphological adaptation. MorphServe dynamically adjusts model precision through LayerSwapper and KVC memory capacity through KVResizer, in a coordinated manner based on real-time resource usage. MorphServe maintains high-quality inference under normal conditions and adapts gracefully during overload periods. Our design achieves substantial improvements in SLO compliance rates, memory efficiency, and serving robustness, while incurring minimal quality loss and runtime overhead.
@@
  • Section A: Related Work
  • Section B: Layer Swapping Sequence Profiling
@@
A  Related Work
LLM Serving Systems. TorchServe [1] and NVIDIA Triton [50] offer general-purpose, modular serving frameworks that enable efficient and parallel inference. Building on these foundations, a wave of systems has emerged to specifically optimize for LLM serving [49, 36, 68, 79, 75, 33, 2, 69, 29]. Orca [75] introduces continuous batching for request-level scheduling to improve throughput. vLLM [33] proposes PagedAttention for fine-grained KV cache management, enhancing memory efficiency at scale. SwiftLLM [58] replicates the core scheduling and memory techniques of vLLM tailored for research, offering a lightweight and extensible platform that retains comparable inference performance. Sarathi [2] introduces a chunked-prefill mechanism that splits requests into smaller segments and piggybacks decoding to better utilize compute resources. FastServe [69] employs iteration-level preemptive scheduling to reduce queuing delays from long-running jobs. While these systems advance scheduling and memory management techniques, they all operate under fixed-precision model assumptions and rely on static memory provisioning at runtime. This limits their ability to respond effectively to volatile or bursty workloads, where demand fluctuates rapidly. In contrast, MorphServe introduces runtime morphological adaptation that enables elastic resource usage tailored to current system load, improving responsiveness while preserving model quality under pressure.
LLM Post-Training Quantization. Post-training quantization (PTQ) is a widely adopted technique for reducing memory and compute costs in LLM serving [13, 32, 34, 76, 38, 17, 71, 37]. Classic round-to-nearest methods [73, 47] offer simplicity but suffer from large quantization errors in outlier-heavy layers. To address this, non-uniform schemes like AdaRound [46] and ZeroQuant [74] introduce pseudo-data–guided codebooks and adaptive rounding, respectively. Calibration-based methods further improve weight quantization: GPTQ [17] uses Hessian-guided optimization for 3–4 bit weights with minimal perplexity loss, while AWQ [38] preserves critical channels using activation statistics to enable accurate 4-bit inference. Recent work extends PTQ to activations to support full matrix-level quantization. SmoothQuant [71] mitigates activation outliers via weight migration for INT8 execution, QServe [39] co-designs weight-activation-KVC with system-level optimizations, and DuQuant [37] uses block-wise rotation and permutation to robustly smooth both normal and massive activation outliers. Despite their effectiveness, all these methods apply static quantization, causing permanent accuracy degradation, even when resources permit full-precision serving. In contrast, MorphServe introduces runtime-selective quantization by dynamically interpolating between pre-generated precision profiles based on system pressure, while also resizing KV cache blocks to handle bursty workloads. Moreover, MorphServe is fully compatible with modern PTQ methods like GPTQ and AWQ, offering an elastic and forward-compatible serving framework.
Mixed-Precision Inference. Recent research has explored mixed-precision execution to improve ML inference efficiency by applying lower-bit computations [54, 28, 25, 66, 8, 18, 14, 55]. Mix-GEMM [54] and Bit-Split [66] focus on hardware-level support for efficient mixed-bit matrix multiplication for DNN inference. PMPD [8] performs prompt-adaptive quantization based on attention entropy. MARLIN [18] proposes a training-free mechanism to generate multiple model variants with differing precision configurations to accelerate end-to-end LLM inference. These approaches either tailor precision to input tokens or statically select precision-aware submodels. While effective in improving FLOPs or latency, they do not consider runtime system factors such as memory saturation, queueing delay, or workload bursts. In contrast, MorphServe introduces the first workload-aware, dynamic mixed-precision serving, where the model’s precision configuration is adjusted dynamically based on real-time system pressure, which enables token-level adaptation that is tightly coupled with runtime memory availability and latency constraints, achieving smoother efficiency–accuracy tradeoffs.
B  Layer Swapping Sequence Profiling
@@
B.1  Layer Importance Score (LIS)
@@
B.2  Greedy Selection Policy

### v2

@@ 4.4 KVResizer: Elastic KVC Resizing
[…]
@@
KVResizer extends PagedAttention Kwon et al. (2023) with kernel-level support for on-demand KVC block allocation/deallocation, implemented through memory mapping without requiring kernel recompilation. All resizing operations are executed asynchronously using separate CUDA streams to avoid interference with ongoing decoding.
Unlike static preallocation strategies (e.g., in vLLM Kwon et al. (2023)), KVResizer adjusts KVC capacity dynamically based on real-time memory availability. Once the pressure subsides, both temporary KVC blocks and quantized layers are released and restored to their full-precision state, ensuring memory reuse and accuracy recovery.
@@ 5 Experiment
Table 1: Impact of mixed-precision serving on Llama 3 8B (BookSum Chapters dataset, 6K input / 2K output tokens).
Generation Strategy F1 Score ROUGE-L Within Accuracy Bound
All tokens with INT4 (W4) 14.4716 12.1598 Lower Bound
First 1K tokens FP16, Last 1K tokens W4 16.6112 15.2346 Yes
First 1K tokens W4, Last 1K tokens FP16 14.7101 12.3213 Yes
First 512 FP16, Middle 1K W4, Last 512 FP16 16.0802 14.4013 Yes
First 512 W4, Middle 1K FP16, Last 512 W4 16.0413 13.1127 Yes
Switch precision every 256 tokens, starting with W4 15.2634 13.0353 Yes
Switch precision every 256 tokens, starting with FP16 16.0012 14.0146 Yes
All tokens with full precision (FP16) 17.6458 15.4847 Upper Bound
Figure 6: MorphServe delays saturation and achieves up to $1.83\times$ throughput than full-precision serving under increasing request rates.
Figure 7: MorphServe incurs negligible runtime overhead while improving tail TPOT latency. MorphServe (green) achieves comparable average TPOT latency to the full-precision baseline (blue), while reducing P99 latency by up to $1.23\times$. Performance mode (light green) improves the average TPOT by up to $1.17\times$ through aggressive layer morphing.
Table 2: Performance of MorphServe with AWQ and uniform INT4 quantization vs. FP16 on Llama 2 7B, using the DuReader dataset and BurstGPT trace. MorphServe improves accuracy over static quantization, reducing accuracy degradation by 83.57% for AWQ, 74.67% for Uniform, while lowering P95 TTFT by over 77% compared to FP16.
Quantization (INT4) Metric Static Quantization MorphServe Full Precision
TTFT P95 (s) 1.1686 1.2420 5.5223
AWQ TPOT P99 (s) 0.0735 0.1064 0.1241
F1 Score 25.55 27.33 27.68
TTFT P95 (s) 1.1321 1.2144 5.5223
Uniform TPOT P99 (s) 0.0728 0.1053 0.1241
F1 Score 24.87 26.94 27.68
Table 3: Cross-dataset evaluation of LIS profiles—perplexity on C4 using LIS calibrated on WikiText-2.
Method 1 2 4 8 16 32 (Fully Quantized)
Front-to-Back 7.0793 7.0822 7.0933 7.1119 7.1524 7.2463
Back-to-Front 7.0871 7.0938 7.1031 7.1224 7.1712 7.2463
Random 7.0871 7.0913 7.1012 7.1234 7.1676 7.2463
LIS 7.0801 7.0818 7.0912 7.1107 7.1497 7.2463
[…]
@@
Implementation. MorphServe is implemented on top of SwiftLLM Jiang et al. (2024); shengyu Liu and Jover Qian , a lightweight and modular LLM inference framework that reproduces vLLM Kwon et al. (2023) performance with simplified components. We added approximately 2,200 lines of Python and 500 lines of C++/CUDA to support MorphServe’s optimized KVC management and attention kernel extensions, which enable efficient layer swapping and KVC resizing at runtime.
Comparison Scope and Baselines. LLM serving frameworks differ widely in objectives: static systems like vLLM Kwon et al. (2023), Sarathi Agrawal et al. (2024), and FastServe Wu et al. (2023) focus on fixed-precision serving, while speculative or input-aware approaches like MARLIN Frantar et al. (2025), LayerSkip Elhoushi et al. (2024), and FlexiDepth Luo et al. (2025) optimize inference paths offline. In contrast, MorphServe adapts in real time, dynamically navigating the performance–accuracy tradeoff during live serving. We benchmark against two representative baselines: FP16 (upper-bound accuracy) and INT4 (performance-optimized static quantization). MorphServe is evaluated in two runtime modes: accuracy mode, which conservatively triggers layer swapping to preserve quality, and performance mode, which aggressively swaps layers to improve throughput under memory pressure. All setups use the same inference engine, and AWQ Lin et al. (2024b) is adopted for its efficient kernel support, though MorphServe is quantization-method agnostic.
@@ 5.1 Main Results
TTFT and Accuracy. As shown in Figure 4, MorphServe significantly reduces P95 TTFT latency (95th-percentile TTFT latency) while preserving output quality in all model-trace-dataset configurations. Compared to full-precision baselines, MorphServe reduces the P95 TTFT by $2.9\times$$15.7\times$ ($2.2\times$$3.9\times$ in accuracy mode and $3.4\times$$19.5\times$ in performance mode) while maintaining quality within $0.51\%$$3.82\%$ degradation on F1 or Rouge-L scores, as low as $0.11\%$$2.18\%$ in accuracy mode. In contrast, static quantization exhibits $2.34\%$-$9.47\%$ degradation compared to full-precision inference, due to suffering from persistent quality loss across the entire serving lifetime. In particular, MorphServe excels in long-context datasets such as GovReport, leveraging LayerSwapper and KVResizer to optimize memory and computing efficiency. MorphServe with different configurations (green stars) visualizes the ability to navigate the latency-accuracy Pareto frontier, offering the best balance of performance and quality based on real-time workload shifts.
Generation Quality under Mixed-Precision Serving. MorphServe adopts a conservative adaptation strategy, replacing full-precision layers with high-quality quantized counterparts (e.g., AWQ) rather than skipping them at the token level et al. . This design preserves stable and bounded generation quality during dynamic adaptation. As shown in Table 1, even for long-context tasks (BookSum Chapters, 6K input and 2K output tokens), all mixed-precision configurations maintain performance between the fully quantized and full-precision bounds, ensuring controllable quality under frequent precision switching. Additional perplexity evaluations across models are provided in Appendix. A verbatim example of MorphServe’s output under mixed-precision serving is also in the Appendix.
@@
Workload Adaptation and Saturation Resilience. As shown in Figure 5, MorphServe adaptively manages KVC block capacity in response to fluctuating load. In the full-precision baseline, KVC usage saturates the static capacity during peak periods, resulting in elevated queueing delays, request preemption, and frequent KVC swapping, which can lead to SLO violations. Static quantization, while reducing the memory footprint, degrades model accuracy and underutilizes GPU memory, even during low-load periods. MorphServe attaches new blocks during bursty traffic and releases them as load subsides, enabled by the synergistic LayerSwapper and KVResizer mechanism. MorphServe improves overall KVC memory utilization and output accuracy by $29.29\%$ and $3.58\%$, respectively, compared to static quantization. The adaptation allows MorphServe to expand KVC usage by up to $32.97\%$ beyond the full-precision limit when needed, and reduce the queueing delay by up to $3.8\times$. MorphServe also mitigates request preemption and KVC swapping under saturation conditions. This enhances system responsiveness and improves token-level efficiency, contributing to reduced end-to-end request latency.
@@
Throughput. In Figure 6, we compare MorphServe with baselines on DuReader under varying request rates. All configurations maintain low TTFT at low RPS, but as load increases, full-precision inference encounters the threshold, where TTFT spikes abruptly due to memory exhaustion and queueing delays. In contrast, MorphServe consistently pushes back this saturation point, achieving $1.6\times$$1.83\times$ higher throughput than full-precision serving across all evaluated models.
@@
TPOT Tail Latency. Figure 7 presents the cumulative distribution (CDF) of time-per-output-token (TPOT) across two datasets and four models under the Azure LLM trace. MorphServe delivers average TPOT comparable to full-precision serving while improving P95 and P99 tail latency by up to $1.06\times$ and $1.23\times$, respectively. These gains arise from eliminating preemption stalls and avoiding KVC swapping or recomputation, the main sources of long-tail delays. In performance mode, TPOT decreases by $1.11$-$1.17\times$, whereas accuracy mode adds up to $1.06\times$ overhead due to preserving more FP16 layers and applying stricter layer-swapping thresholds, occasionally causing short queuing under load. The improvement results from faster inference on quantized layers and highly efficient swapping kernels, confirming that MorphServe introduces negligible runtime overhead while effectively reducing tail latency.
5.2 Ablation Study
Effectiveness Across Quantization Algorithms. MorphServe is agnostic to specific quantization schemes. We evaluate it under both AWQ and uniform INT4 using identical datasets and traces. As shown in Table 2 (and Table 5 in the appendix), MorphServe consistently improves F1 and Rouge-L over static quantization while maintaining significant latency advantages over full-precision serving. These results confirm that its gains arise from runtime adaptation rather than the choice of quantization algorithm.
Generalization of Layer Importance Score. Table 3 shows the generalization of LIS profiles, which consistently outperforms front-to-back, back-to-front, and random layer orders on C4 with calibration on Wiki2, indicating that LIS captures intrinsic layer importance and generalizes across datasets without re-profiling.
6 Discussion and Future Work
User Tolerance. Since fully quantized models (e.g., AWQ) are widely acceptable, MorphServe’s outputs remain within user tolerance. It maintains full-precision decoding under light load and bounds quality variation under adaptation, achieving stable perceptual quality with adaptive efficiency.
Scalability. MorphServe scales naturally to distributed serving with data, tensor, pipeline, and sequence parallelism. Each GPU runs an independent executor that swaps layers and resizes KV caches locally.
Fairness. The continuous batching Yu et al. (2022) preserves FIFO ordering, while chunked prefill Agrawal et al. (2024) and fine-grained KV management Kwon et al. (2023) prevent starvation. Combined with these, MorphServe serving ensures fair and stable serving under bursty conditions.
Future Work. MorphServe can extend beyond weight-only quantization to include activation and KV quantization. Such extensions promise higher efficiency but raise calibration and scheduling challenges. We plan to explore these for edge deployment and robust, energy-efficient adaptation.
@@ 7 Conclusion
This paper presents MorphServe, a novel workload-aware LLM serving framework based on morphological adaptation. Guided by a unified Layer Importance Score, MorphServe dynamically adjusts model precision through LayerSwapper and KV cache memory capacity through KVResizer, in a coordinated manner based on real-time resource usage. MorphServe maintains high-quality inference under normal conditions and adapts gracefully during overload periods. Extensive experiments show that MorphServe improves SLO compliance, memory efficiency, and robustness, while maintaining high generation quality along the efficiency–accuracy Pareto frontier.
@@
  • Section A: Layer Swapping Sequence Profiling
  • Section B: Additional Experimental Results
@@
A  Layer Swapping Sequence Profiling
@@
A.1  Layer Importance Score (LIS)
@@
Hyperparameter Selection. To balance robustness and accuracy, we define the hyperparameters of LIS as follows: $\alpha_{1}$ captures weight sensitivity (e.g., norm-based changes after quantization), $\alpha_{2}$ measures quantization distortion in activation space (e.g., cosine similarity), and $\beta$ reflects model-level degradation (e.g., perplexity under greedy layer removal).
Table 4: Perplexity comparison of LIS under different ($\alpha_{1}$$\alpha_{2}$$\beta$) hyperparameter settings across quantization levels on Llama 2 7B using WikiText-2.
# Layers Quantized 1 2 4 8 16 32 (Fully Quantized)
LIS (0.33,0.33,0.33) 5.4732 5.4748 5.4786 5.4905 5.5257 5.6002
LIS (0.25,0.25,0.5) 5.4732 5.4743 5.4779 5.4875 5.5215 5.6002
We empirically choose $\alpha_{1}=\alpha_{2}=0.25$ and $\beta=0.5$, ensuring $\alpha_{1}+\alpha_{2}+\beta=1.0$. This heuristic gives slightly more emphasis to $\beta$ due to its stability across inputs and its direct alignment with user-perceived output quality. Relying only on $\beta$, however, risks overlooking layers with high local distortion, while $\alpha_{1}$ and $\alpha_{2}$ alone are more sensitive to input variance. We compare this configuration to a baseline with uniform weighting ($\alpha_{1}=\alpha_{2}=\beta=0.33$). As shown in the Table 4, our setting produces lower perplexity across quantization levels and leads to more stable LIS rankings.
@@
Table 5: Perplexity comparison of LIS using L2 Norm vs. cosine similarity across quantization levels on Llama 2 7B with WikiText-2.
# Layers Quantized 1 2 4 8 16 32 (Fully Quantized)
LIS - L2 Norm 5.4733 5.4776 5.4848 5.5021 5.5321 5.6002
LIS - Cosine Similarity 5.4732 5.4743 5.4779 5.4875 5.5215 5.6002
@@
A.2  Greedy Selection Policy